If in the above example, the manager of Universal autos wants the probability of the stockout to be no more than 4%, then what should the reorder point be? Assuming that it has been determined that the demand during replenishment lead-time is normally distributed with a mean of 15 gallons and a standard deviation of 6 gallons.


1
Expert's answer
2022-06-01T10:35:59-0400
"P(X<x)=P(Z<\dfrac{x-15}{6})=0.04"

"\dfrac{x-15}{6}=-1.750686"

"x=15-6(1.750686)"

"x=4.496"